the charge on 1 g of Al3+ ions is:
  1. 1/27NA e coulomb
  2. 1/3NAe coulomb
  3.  1/9NA e coulumb
  4.  3NA ecoulomb
 

Grade:11

7 Answers

Dayitva Goel
6 years ago
Charge on one mole (27 g) of Al3+ ions = Ne coulomb
Charge on 1 g of Al3+ ions = 1/27 N?e coulomb
where NA is Avagadro’s number (6.02 × 1023).
Ayushi
6 years ago
1gm atoms =1 mole atoms 1 gm molecules =1mole molecules Same as ...1 gm ion is written in place of 1 mole . i.e 1 gm ion = 1 mole So charge on one mole =3 × NA e-
Mohit
5 years ago
Since we know that GRAM IONS are also called 1mol of ions.So the question here is about charge on 1 mol ions of Al+++.
Now,
No. Of electrons in 1 atom of Al+++. Is 10.(13-3)
Charge in 1 electron is 1.6*10^-19.
So, charge in 1atom of Al+++ is 10*1.6*10^-19
Therefore,charge in 1 mol ions = (16*10^-19)*avogadro no.
Will give 3*avogadro e coulomb.
Charge in one e
Zeeshan bashir
5 years ago
Firsty no. Of atoms(N) =no. Of moles(n)×Na
n =1/27
N=1/27×Na since one atom contains 3e therefore N no. Of atoms contain N×3e
so charge on 1 g al3+= 1/27 ×3e × Na = 1/9 × e columb× Na
Kushagra Madhukar
3 years ago
Hello Student
 
The charge on 1 ion of Al3+ = 3 eC
Hence,
Charge on 1 mole of Al3+ ions = 3 x NA eC
27 g of Al3+ ions = 3 x NA eC (molar mass of Al = 27)
1 g of Al3+ ions = 3/27 NA ec = 1/9 NA eC
 
Hope it helps
Yash Chourasiya
3 years ago
Dear Student

The charge on one ion of Al3+= 3 eC
Hence,
Charge on 1 mole of Al3+ions = 3 x NAeC
27 g of Al3+ions = 3 x NAeC (molar mass of Al = 27)
1 g of Al3+ions = 3/27 NAeC
= 1/9 NAeC

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya
Singari Sai Ashwatha
2 years ago
The charge on 1 ion of Al3+ = 3 eC
27 g of Al3+ ions = 3 x NA eC
1 g of Al3+ ions = 3/27 NA ec = 1/9 NA eC